# AMBER: Algorithm for Multiplexing spectrometer Background Estimation with Rotation-independence

### v1

@@ Abstract
C1 Current code version 1.0.0
@@ 2.2 Mathematical Formulation
$\min_{X,b}$ $\frac{1}{2}\lVert Y-X-\mathcal{R}b\rVert_{2}^{2}+\lambda||X||_{1}+\frac{\beta}{2}\mathrm{Tr}\left(b^{T}L_{b}b\right)+\frac{\mu}{2}\boldsymbol{1}_{n_{x}}^{T}X^{T}L_{\omega}X\boldsymbol{1}_{n_{y}}$ (1)
@@ 2.3 Coordinate descent algorithm to solve the minimization problem
The loss function of Eq. 1 for a given signal $X$ and background $b$ is denoted by $\mathcal{L}(X,b)$ Equation. 1 can in principle be solved using proximal gradient based algorithm but a significant number of iterations may be required to reach an optimal solution. Instead, we use a coordinate descent algorithm, in which $\mathcal{L}$ is iteratively minimized with respect to $X$ keeping $b$ fixed and with respect to $b$ keeping $X$ fixed. This method enables to use closed-form updates. Given initial values of the signal $X^{0}$ and the background and $b^{0}$, we iterate $K$ times for $k=0,\ldots,K-1$,
@@
where $\Gamma$ is a $n_{q}\times n_{q}$ diagonal matrix whose elements are the number of measurements in each radial bin, $\mathcal{R}^{*}$ is the adjoint of operator $\mathcal{R}$ (i.e., $\langle\mathcal{R}U,V\rangle=\langle U,\mathcal{R}^{*}V\rangle$ where $\langle$,$\rangle$ , ⟩ is the standard inner product in Euclidean space) and $S_{u}(x)=\mathrm{sign}(x)\max(|x|-u,0)$. One can show that $\mathcal{R}^{*}$ maps any $n_{x}\times n_{y}\times n_{\omega}$ matrix into a $n_{q}\times n_{\omega}$ matrix, whose elements are the circular sum across angles for each radial bin (see B for the detailed proof.) The main advantage of such an approach is that both subproblems (Eqs.  (4) and  (5)) have closed-form solutions. One can observe that the inverse of the matrices $\boldsymbol{I}+\mu L_{\omega}$ and $\Gamma+\beta L_{q}$ do not depend on the data and they can be precomputed ahead of time.
2.4 Hyper-parameter tuning
@@ 2.4.1 Heuristic Parameter Determination
    $\beta$: This parameter corresponds to a smoothing penalty on the background. It can be estimated through a cross-validation procedure including two steps. Data points containing mainly signal are removed through a thresholding, leaving points assumed to contain mostly background, i.e. $X\approx 0$. This results in $B\sim Y$ and thus one uses the RMS of $Y-B$ as a measure. The thresholding consists of masking values that are above a given quantile, $q$, of the observed values $Y$.
The approach provides a generic initial setting of the hyperparameters. However, the strategy can be suboptimal depending on the acquired measurements and the sample. It is also mentioned that the determination of $\beta$ is potentially time consuming depending on the range of possible values. Hence, hyperparameter tuning can be improved based on instrument and spectrum knowledge (smoothness, sparsity, etc.). However, the computation time of our denoising algorithm is low - on the order of 1 minute - allowing the user to evaluate the signal decomposition for a wide range of hyperparameters.
@@ 2.5 Code implementation
As an example, the latest MJOLNIR[16] version (1.3.5) contains a dedicate background object through which the AMBER decomposition algorithm can be performed. MJOLNIR has been written for a wide range multiplexing spectrometers and can thus be used in many neutron scattering experiments. In particular, data is binned along $Q_{x}$, $Q_{y}$ and $\omega$. The decomposition process is semi-automatised with standard values for the hyperparameters, but some tuning may be required. Firstly, the binning parameters $n_{x},n_{y},n_{\omega}$ and $n_{q}$ are chosen. Secondly, the algorithm hyper-parameters $(\lambda,\beta,\mu)$ can be either set manually based on expert knowledge or using the heuristic approach mentioned in the sec. 2.4.1.
@@ 2.6 Limitations
The elastic line - the entire scattering plane at 0 energy and energy transfer within the instrument resolution - is a natural region in which all assumptions are violated. More specifically, assumption 1 is invalid as sample holders usually have texture or a non-isotropically shape. The abrupt change of intensity between elastic and inelastic scattering invalidates assumption 2 while Bragg peaks invalidates the smoothness and continuity of the signal in energy and $|\vec{Q}|$. Accidental Bragg scattering, denoted as Currat-Axe Spurions, [3, 9] also violate the assumptions, in particular assumption 3. Lastly, for samples where $S(\vec{Q},\omega)$ is dominated by a constant, scattering vector-independent signal, e.g. potentially the top of a magnon band or crystal electric fields, challenges assumption 3.
@@ 3 Numerical experiment
The data were acquired using CAMEA with instrument settings tabulated in Tab. 2. The sample was measured within the horizontal ($H$,0,$L$) plane. The data reveal a low signal-to-noise level and a complex excitation spectrum shown in Fig. Fig. 2(a). It contains a series of spin waves extending up to 6.5 meV. While the magnetic excitations disperse mainly along the crystallographic $H$ direction, the $L$ dependence is limited. Thus, the data is presented in cuts along $H$ and energy transfer. The four strongest sample signatures are the three spin waves dispersing between 1 and 1.6 meV and 2.2 and 3 meV, and a less dispersive one around 3.7 meV. The lowest mode is visible across the entire $H$ range, revealing minima at integer values of $H$. The second magnon features equal minima, but has a modulated intensity. The higher mode is only visible around odd values of $H$, and is superimposed with a flat mode at 3.7 meV. Only data above the elastic line, i.e. $\omega$ $>$ 0.5 meV and without spurions are shown.
@@
We compare the performance of AMBER against an expert background estimation, stemming from a manual mask of the foreground signals. This roughly 8 hours work allows to find the background signal of each data file by averaging along constant $|\vec{Q}|$ and is subtracted from the data. This is a common method already used in other studies[21, 22]. The results for VOSe2O5 are shown in Fig. 2(c). AMBER was run with a binning $Q_{x}=Q_{y}=$0.03 Å-1 and $E=$0.05 meV. $\lambda$ and $\mu$ were found according to sec. 2.4.1, and $\beta$ through the cross validation method with a background threshold of 60 %. The final RMSE for this cross validation is 107.98, which required a validation of 28 iterations resulting in a computational time of 17 min (each iteration took $\sim$ 50 s). The results presented in Fig. 2(b) were obtained with the parameters shown in Tab. 3.
@@
Figure 2: Scattering intensities of VOSe2O5 as a function of energy transfer and scattering vector along ($H$,0,1) showing (a) the raw data, (b) subtracted data using AMBER and (c) the results from an expert performing a manual subtraction. All plots were performed using an integration of 0.56 Å orthogonal to the cut, 0.014 Å along the cut, and 0.05 meV along the energy axis.
@@ 4 Impact
The data collection strategy of modern neutron spectrometers such as the multiplexing instrument CAMEA allow for a simultaneous measurement of foreground and background features. The current best-practice methodology for background determination is to manually mask areas containing foreground features, and use the remaining data as background. An interpolation among background points enables an estimation of the background below the signal taking advantage of the sample rotation independence (see Sec. 2.1). However, this process is labour intensive and requires expert-knowledge about the expected signal and about the instrument. The process is also highly volatile to subjective errors. AMBER allows non-expert users to perform an efficient foreground-background segmentation of three-dimensional data sets obtained on multiplexing spectrometer. Using AMBER  the background determination is reduced to minutes, it is objective and reproducible. This is especially practical for measurements suffering from low signal-to-noise ratios. However, we advice users not to blindly trust the algorithm output.
A logical expansion path of AMBER is to include time-of-flight neutron spectroscopy data, which is also based on rotation scans following the same assumptions as described in sec. 2.1. Since the standard time-of-flight analysis software, Mantid[23], is build on Pyhton, we are positive that an implementation of AMBER is achievable with relatively low efforts. AMBER may also be beneficial for x-ray and electron based experimental techniques, such as angle-resolved photoemission, resonant inelastic x-ray scattering, tunnelling electron spectroscopy and other techniques using large area detectors.
@@ A.1 Synthetic data - MnF2
[…]
@@
[…]
@@ A.2 Experimental data - MnF2
The data are visualized in Fig. 10 showing scattering intensity as a function of scattering vector along $(-1,0,L)$. The figure is split into three panels showcasing (a) the ground truth, i.e. using the standard background masking method (b) the experimental data, (c) background subtracted data utilizing AMBER. In the case of MnF2, the masking method removes signal around predicted spin-wave positions. The three figures are very similar, with slight differences highlighted through 1D cuts along constant energy and constant $|Q|$. These are presented in Fig. 11. Figures (a) and (c) show the scattering intensity overplotted with the estimated background from the two methods, and (b) and (d) show the resulting background-subtracted intensities. For the constant energy cut, AMBER and the masking method perform very similarly with barely any difference while for constant $|Q|$ AMBER performs a slight over-subtraction at higher energies.
@@ Appendix C Energy-dependent parameters
$\min_{X,b}$ $\frac{1}{2}\lVert Y-X-\mathcal{R}b\rVert_{2}^{2}+\sum_{e=1}^{n_{\omega}}\boldsymbol{\lambda}||X_{e}||_{1}+\sum_{e=1}^{n_{\omega}}\frac{\boldsymbol{\beta}}{2}\mathrm{Tr}\left(b_{e}^{T}L_{b}b_{e}\right)+\frac{\mu}{2}\boldsymbol{1}_{n_{x}}^{T}X^{T}L_{\omega}X\boldsymbol{1}_{n_{y}}.$
@@
S1 Current software version 0.0.3

### v2

@@ Abstract
C1 Current code version 1.0.1
@@ 1 Introduction
An alternative to both AMBER and the above described masking method is to prepare the experimental system in a setup where background is directly measured. However, this cannot always be achieved, either due to the sample properties, the lack of availability beam time or the sample environment utilized. However, such backgrounds are expected to be similarly good or better than a background estimate from AMBER. An alternative method often used in ToF experiments is a measurement of an empty sample holder. However, both approaches are time consuming.
@@ 2.2 Mathematical Formulation
[…]
@@ 2.3 Coordinate descent algorithm to solve the minimization problem
The loss function of Eq. 1 for a given signal $X$ and background $b$ is denoted by $\mathcal{L}(X,b)$ . Equation. 1 can in principle be solved using proximal gradient based algorithm but a significant number of iterations may be required to reach an optimal solution. Instead, we use a coordinate descent algorithm, in which $\mathcal{L}$ is iteratively minimized with respect to $X$ keeping $b$ fixed and with respect to $b$ keeping $X$ fixed. This method enables to use closed-form updates. Given initial values of the signal $X^{0}$ and the background and $b^{0}$, we iterate $K$ times for $k=0,\ldots,K-1$,
@@
where $\Gamma$ is a $n_{q}\times n_{q}$ diagonal matrix whose elements are the number of measurements in each radial bin, $\mathcal{R}^{*}$ is the adjoint of operator $\mathcal{R}$ (i.e., $\langle\mathcal{R}U,V\rangle=\langle U,\mathcal{R}^{*}V\rangle$ where $\langle,\rangle$ is the standard inner product in Euclidean space) and $S_{u}(x)=\mathrm{sign}(x)\max(|x|-u,0)$. One can show that $\mathcal{R}^{*}$ maps any $n_{x}\times n_{y}\times n_{\omega}$ matrix into a $n_{q}\times n_{\omega}$ matrix, whose elements are the circular sum across angles for each radial bin (see B for the detailed proof.) The main advantage of such an approach is that both subproblems (Eqs.  (4) and  (5)) have closed-form solutions. One can observe that the inverse of the matrices $\boldsymbol{I}+\mu L_{\omega}$ and $\Gamma+\beta L_{q}$ do not depend on the data and they can be precomputed ahead of time.
2.4 Hyperparameter tuning
@@ 2.4.1 Heuristic Parameter Determination
    $\beta$: This parameter corresponds to a smoothing penalty on the background. It can be estimated through a cross-validation procedure including two steps. Data points containing mainly signal are removed through a thresholding, leaving points assumed to contain mostly background, i.e. $X\approx 0$. This results in $B\sim Y$ and thus one uses the RMS of YBY-B as a measure. The thresholding consists of masking values that are above a given quantile, $q$, of the observed values $Y$.
The approach provides a generic initial setting of the hyperparameters. However, the strategy can be suboptimal depending on the acquired measurements and the sample. It is also mentioned that the determination of $\beta$ is potentially time consuming depending on the range of possible values. Hence, hyperparameter tuning can be improved based on instrument and spectrum knowledge (smoothness, sparsity, etc.). However, the computation time of our denoising algorithm is low - on the order of 1 minute - allowing the user to evaluate the signal decomposition for a wide range of hyperparameters. To check the applicability of AMBER on a given dataset it is suggested to perform the following steps. Firstly, identify the signal to background level, i.e. estimate the percentage of data which contain only background. Next, estimate both $\lambda$ and $\mu$ as described above and find $\beta$ through a cross-validation. This provides a starting point to perform the first background subtraction. In cases where AMBER does not perform satisfactory, as exemplified in 2.6, the subtracted data will have substantial regions of negative intensity or artefacts from abrupt changes in the background estimation. These effects signify cases in which the heuristic to determine the hyperparameters leads to less accurate background estimates. However, it is not excluded that a set of parameters would result in a satisfactory background estimation.
@@ 2.5 Code implementation
As an example, the latest MJOLNIR[16] version (1.3.5) contains a dedicate background object through which the AMBER decomposition algorithm can be performed. MJOLNIR has been written for a wide range multiplexing spectrometers and can thus be used in many neutron scattering experiments. In particular, data is binned along $Q_{x}$, $Q_{y}$ and $\omega$. The decomposition process is semi-automatised with standard values for the hyperparameters, but some tuning may be required. Firstly, the binning parameters $n_{x},n_{y},n_{\omega}$ and $n_{q}$ are chosen. Secondly, the algorithm hyperparameters $(\lambda,\beta,\mu)$ can be either set manually based on expert knowledge or using the heuristic approach mentioned in Sec. 2.4.1.
@@ 2.6 Limitations
The elastic line - the entire scattering plane at 0 energy and energy transfer within the instrument resolution - is a natural region in which all assumptions are violated. More specifically, assumption 1 is invalid as sample holders usually have texture or a non-isotropically shape. The abrupt change of intensity between elastic and inelastic scattering invalidates assumption 2 while Bragg peaks invalidates the smoothness and continuity of the signal in energy and $|\vec{Q}|$. Accidental Bragg scattering, denoted as Currat-Axe Spurions, [3, 9] also violate the assumptions, in particular assumption 3. Lastly, for samples where $S(\vec{Q},\omega)$ is dominated by a constant, scattering vector-independent signal, e.g. potentially the top of a magnon band or crystal electric fields, challenges assumption 3. We mention that in cases of strongly absorbing samples or a highly irregular sample holders, the rotation independence is no longer valid, even deeply inelastically. An absorption correction can be applied prior to the background model while the sample holder case cannot be treated with AMBER.
We further note that AMBER is also applicable for low-dimensional signals, in which the neutron scattering signal along some crystallographic directions remain unchanged. The main requirement for AMBER to perform well is the presence of sufficient background volume at all constant energy planes.
@@ 3 Numerical experiment
The data were acquired using CAMEA with instrument settings tabulated in Tab. 2. The sample was measured within the horizontal (HH,0,$L$) plane. The data reveal a low signal-to-noise level and a complex excitation spectrum shown in Fig. Fig. 2(a). It contains a series of spin waves extending up to 6.5 meV. While the magnetic excitations disperse mainly along the crystallographic $H$ direction, the $L$ dependence is limited. Thus, the data is presented in cuts along $H$ and energy transfer. The four strongest sample signatures are the three spin waves dispersing between 1 and 1.6 meV and 2.2 and 3 meV, and a less dispersive one around 3.7 meV. The lowest mode is visible across the entire $H$ range, revealing minima at integer values of $H$. The second magnon features equal minima, but has a modulated intensity. The higher mode is only visible around odd values of $H$, and is superimposed with a flat mode at 3.7 meV. Only data above the elastic line, i.e. $\omega$ >> 0.5 meV and without spurions are shown.
@@
We compare the performance of AMBER against an expert background estimation, stemming from a manual mask of the foreground signals. This roughly 8 hours work allows to find the background signal of each data file by averaging along constant $|\vec{Q}|$ and is subtracted from the data. This is a common method already used in other studies[21, 22]. The results for VOSe2O5 are shown in Fig. 2(c). AMBER was run with a binning $Q_{x}=Q_{y}=$0.03 Å-1 and E=E=0.05 meV. $\lambda$ and $\mu$ were found according to sec. 2.4.1, and $\beta$ through the cross validation method with a background threshold of 60 %. The final RMSE for this cross validation is 107.98, which required a validation of 28 iterations resulting in a computational time of 17 min (each iteration took $\sim$ 50 s). The results presented in Fig. 2(b) were obtained with the parameters shown in Tab. 3.
@@
Figure 2: Scattering intensities of VOSe2O5 as a function of energy transfer and scattering vector along (HH,0,1) showing (a) the raw data, (b) subtracted data using AMBER and (c) the results from an expert performing a manual subtraction. All plots were performed using an integration of 0.56 Å orthogonal to the cut, 0.014 Å along the cut, and 0.05 meV along the energy axis.
@@ 4 Impact
The data collection strategy of modern neutron spectrometers such as the multiplexing instrument CAMEA allow for a simultaneous measurement of foreground and background features. The current best-practice methodology for background determination in already measured data is to manually mask areas containing foreground features, and use the remaining data as background. An interpolation among background points enables an estimation of the background below the signal taking advantage of the sample rotation independence (see Sec. 2.1). However, this process is labour intensive and requires expert-knowledge about the expected signal and about the instrument. The process is also highly volatile to subjective errors. AMBER allows non-expert users to perform an efficient foreground-background segmentation of three-dimensional data sets obtained on multiplexing spectrometer. Using AMBER  the background determination is reduced to minutes, it is objective and reproducible. This is especially practical for measurements suffering from low signal-to-noise ratios. However, we advice users not to blindly trust the algorithm output.
A logical expansion path of AMBER is to include time-of-flight neutron spectroscopy data, which is also based on rotation scans following the same assumptions as described in sec. 2.1. Since the standard time-of-flight analysis software, Mantid[23], is build on Python, we are positive that an implementation of AMBER is achievable with relatively low efforts. AMBER may also be beneficial for x-ray and electron based experimental techniques, such as angle-resolved photoemission, resonant inelastic x-ray scattering, tunnelling electron spectroscopy and other techniques using large area detectors. The mathematical formulation behind the algorithm extends to a third dimension of $\vec{Q}$ either through an expansion of the found background estimate from 2 to 3D, i.e. being dependent of $|\vec{Q}|_{x,y}$ calculated from the in-plane scattering vector, the out of plane scattering vector and energy. Alternatively, one could include $\vec{Q}_{z}$ directly in the calculation of $|\vec{Q}|$ or ignore it altogether. Both of these suggestions impose assumptions not based on known physical background behaviour. A far greater issue is that of computational complexity. The computational complexity of one iteration step would significantly increase due to the matrix inversion. Alternatively, if the algorithm is made tractable it could be distributed across a multi-kernel setup like a cluster or a graphical unit.
@@ A.1 Synthetic data - MnF2
We evaluated the performance of AMBER against the baseline for 10 values of $\gamma$ equally spaced in $1.0\mathrm{e}2;1.0\mathrm{e}3]$ and we fix the amplitude of the background $A=100.0$.
@@
Figure 8: RMSE of AMBER (blue) and the baseline approach (red) with respect to Signal-to-Noise ratio for each value of $\gamma$ in $1.0\mathrm{e}2;1.0\mathrm{e}3]$.
@@ A.2 Experimental data - MnF2
The data are visualized in Fig. 10 showing scattering intensity as a function of scattering vector along (1,0,L)(-1,0,L). The figure is split into three panels showcasing (a) the ground truth, i.e. using the standard background masking method (b) the experimental data, (c) background subtracted data utilizing AMBER. In the case of MnF2, the masking method removes signal around predicted spin-wave positions. The three figures are very similar, with slight differences highlighted through 1D cuts along constant energy and constant $|Q|$. These are presented in Fig. 11. Figures (a) and (c) show the scattering intensity overplotted with the estimated background from the two methods, and (b) and (d) show the resulting background-subtracted intensities. For the constant energy cut, AMBER and the masking method perform very similarly with barely any difference while for constant $|Q|$ AMBER performs a slight over-subtraction at higher energies.
@@ Appendix C Energy-dependent parameters
[…]
@@
S1 Current software version 1.0.1